Circular No. 8961

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

η CARINAE

J. C. Martin, University of Illinois at Springfield; K. Davidson, University of Minnesota; and collaborators (R. M. Humphreys and A. Mehner, University of Minnesota; F. Hamann, University of Florida; G. Ferland, University of Kentucky; and K. Ishibashi, Northwest Research Associates) report the first definite sign of η Car’s next spectroscopic event that is predicted to occur in 2009 January. Broad He II 468.6-nm emission has appeared in blue and red Gemini/GMOS spectra obtained on 2008 July 4 and 17. The central velocity is near −40 km/s heliocentric, the FWHM is roughly 300 km/s, and the equivalent width is ∼ 0.01 nm. In η Car (HD 93308), this very-high-excitation feature becomes detectable only near a spectroscopic event, and it did not appear in similar data obtained during 2007 June-July and 2008 February. No existing model clearly predicts its appearance as early as six months before the oncoming event (Martin 2006, Ap.J. 640, 474, and references therein). The only other perceptible spectral changes at this time involve the He I triplet (but not singlet) features. P-Cyg absorption around −400 km/s in the He I 402.6-, 447.2-, and 471.3-nm emission lines has increased by ∼ 50 percent since mid-2007. Such absorption is expected to occur in outer regions of the primary wind photoionized by the hot secondary star, if the latter is now moving approximately between us and the primary. Thus it helps to constrain the orbit orientation. For background information see Ishibashi 2001, ASP Conf. Proc. 242, 53; Nielsen et al. 2007, Ap.J. 660, 669; Humphreys et al. 2008, A.J. 135, 1249;

and Okazaki et al. 2008, http://lanl.arxiv.org/abs/0805.1794v1.

COMETS C/2008 H6–H9 (SOHO) Further to IAUC 8957 and 8960, additional Kreutz sungrazers have

been found on SOHO website images. C/2008 H6 was very faint (mag ∼ 8) and diffuse. C/2008 H7 was very diffuse (mag ∼ 7.5–8) with a faint tail. C/2008 H8 was extremely faint (mag ∼ 8–8.5) and slightly diffuse. C/2008 H9 was tiny and stellar in appearance (mag ∼ 8).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 H6 Apr. 27.188 2^{h}25.^{m}7 +12^{o}46′ C2 EB 2008-L47
C/2008 H7  29.830 2 35.4 +13 31 C2 MM 2008-L47
C/2008 H8  30.881 2 39.1 +13 50 C2 MK 2008-L47
C/2008 H9  21.688 2 06.3 +11 05 C2 HS 2008-M10
2008 July 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green